OpenStudy (nathalyn):

HHHHHHHHHHHHHHHHEEEEEEEEEEEEEEEEEEEEEEEEEELLLLLLLLLLLLLLLLLLLLLLLLPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPMMMMMMMMMMMMMMMMMMMMMEEEEEEEEEEEEEEEEEEEEEEEEEEEEEIIIIIIIIIIITTTTTTTTTTTTTTSSSSSSSSSUUUUUUURRRRRRRRGGGGGGGEEEEEEENNNNNNNNNNNNNTTTTTTTTTTTT!!!!!!!!!!!!!!!!!!!! 3+ - 18t + 4 = -26

pooja195 (pooja195):

Combine like terms (−18t)+(3+4)=−26

OpenStudy (nathalyn):

but I need the whole process thats what I can do!! :(((

pooja195 (pooja195):

3−18t+4=−26 3+−18t+4=−26 (−18t)+(3+4)=−26(Combine Like Terms)

pooja195 (pooja195):

3+4=?

OpenStudy (nathalyn):

um..

pooja195 (pooja195):

What im showing u the steps?

OpenStudy (nathalyn):

s3+4=?? ( i need to give the answer right?/

pooja195 (pooja195):

3−18t+4=−26 3+−18t+4=−26 (−18t)+(3+4)=−26(Combine Like Terms) −18t+7=−26 −18t+7=−26 Subtract 7 from both sides. −18t+7−7=−26−7 −18t=−33 Divide both sides by -18 -18/-18=t -33/-18=?

OpenStudy (nathalyn):

-18?

pooja195 (pooja195):

-33/-18=?

OpenStudy (nathalyn):

I mean the answer is -18 ??

pooja195 (pooja195):

You cant divide -33/-18 so it would be a fraction you end up with t=11/6

OpenStudy (nathalyn):

okay so is -1

pooja195 (pooja195):

t=11/6 is the final answer .-. its already simplified

OpenStudy (nathalyn):

oh okay thank very much i understad know thnks!!!!!

pooja195 (pooja195):

:)